The vertices of a triangle A(-1,-7) B(5,1) C(1,4) the equation of angle bisector of  angle ABC

Grade:11

3 Answers

Arun
5 years ago
If the bisector of angle B meets side AC at D then 
AD/DC = BA/BC. 
BA^2=(5--1)^2+(1--7)^2=36+64=100 
BA=10 
BC=(5-1)^2+(1-4)^2=16+9=25 
BC=5 
so 
AD/DC=10/5=2. 

The co-ordinates of D are  
(2/3)C+(1/3)A=(2/3)(1,4)+(1/3)(-1,-7) 
=(1/3)(1,1) 
=(1/3,1/3). 

The slope of DB is m=(1-1/3)/(5-1/3)=(2/3)/(14/3)=1/7. 
The line BD passes through B so it has equation 
y-1=(1/7)(x-5) 
7y-7=x-5 
7y=x+2. 

ALTERNATIVELY: 
The slope of AB is m1=(1--7)/(5--1)=8/6=4/3, the slope of BC is m2=(4-1)/(1-5)= -3/4. m1m2=-1 so AB is perpendicular to BC hence angle B is a right angle. If the slope of BC is m2=tanθ= -3/4 then the slope of the bisector of angle B is  
tan(θ+45º)=(tanθ+tan45º)/(1-tanθtan45º) 
= (-3/4+1)/(1-(-3/4)*1) 
= (1/4)/(7/4) 
= 1/7. 
The bisector passes through B(5,1) so its equation is 
y-1=(1/7)(x-5) 
7y=x+2.
Rajdeep
5 years ago
HELLO THERE!
 
The process to solve such problems is that, first find the angle between the two lines. The angle bisector will pass through a given point (here, point B). And, when we get the angle between the two lines, we can find out the slope of the angle bisector, as the angle that the angle bisector will make with one of the straight lines will be half of the angle between the two linesFrom the slope, and the point through which the angle bisector bisects, we can find out the equation of the angle bisector.
 
LET’S GO!
 
We have A(-1, -7)
B(5, 1) and C(1, 4).
 
Slope of line AB:
\frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \frac{1+7}{5+1} = \frac{4}{3} = m_{1}
 
Slope of line CB:
\frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \frac{1-4}{5-1} = -\frac{3}{4} = m_{2}
 
Now, to find the angle between the two lines whose slopes are known, we have the formula:
 
tan \theta = |\frac{m_{1}-m_{2}}{1+m_{1}m_{2}}| \\\\\implies tan\theta = |\frac{{}\frac{4}{3}+\frac{3}{4}}{1-1}| \\\\\implies tan\theta = undefined \\\\\implies \theta = 90\degree
 
Now, when angle between the two straight lines is 90 degrees, the angle between the angle bisector through B and line AB is 45 degrees. So, slope of angle bisector = tan 45 = 1.
 
Now, from the equation:
y - y_{1} = m(x - x_{1}) \\\\\implies y - 1 = 1(x - 5) \\\\\implies x - y = 4
 
This is the equation of the angle bisector. The angle bisector goes through the point B (5, 1) and its slope is 1, so we found out its equation by the above process. THANKS!
Rajdeep
5 years ago
Sorry, I have committed a mistake in the above answer.
Actually the angle bisector makes the angle 45 degrees with the negative X-axis.
So, with the positive X axis it makes 180 – 45 = 135 degrees.
 
So its slope becomes tan 135 = -1.
 
The equation becomes y – 1 = -1 (x – 5)
=> x + y = -4
 
THANKS!